OpenStudy (anonymous):

Which of these expressions resolves to a finite set of solutions for x? a.)|2 + x| = 7 b.)2 ≤ x ≤ 7 c.)|2 + x| > 7 d.)2 < x < 7

OpenStudy (nowhereman):

Please find the solutions for (a).

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

please help

OpenStudy (nowhereman):

Which values for x make the equation in a.) true?

OpenStudy (anonymous):

2 & 7

OpenStudy (nowhereman):

No, that is not true. I'll give you a definition: \[|a| = \begin{cases}a\quad \text{if}\;a \geq 0\\-a\quad\text{if}\;a < 0\end{cases}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think the answer is A

OpenStudy (anonymous):

Because x can only be 5 or -9 if |x+2|=7